bejaminmiller:

i need help total distance 385 plane distance 346 What percentage of the total distance did Kiley travel by plane?

martai:

what do u think it is ?

bejaminmiller:

i dont know

martai:

can u tell me what are the answers?

bejaminmiller:

its a question ware you type the answers

martai:

oh:/

bejaminmiller:

can you help?

surjithayer:

\[\frac{ 346 }{ 385 }\times100\approx 89.87 ~\% \]

bejaminmiller:

thank you!

surjithayer:

yw

Alexis1415:

@surjithayer wrote:
\[\frac{ 346 }{ 385 }\times100\approx 89.87 ~\% \]
Could you explain how you did this process and how you came up with that?

surjithayer:

out of total distance traveled 385, distance traveled by plane=346 " " " " " 1 , " " " " =346/385 " " " " " 100, " " " " \[\frac{346}{385}\times 100 \approx 89.87 ~\%\]